To test whether all University students spent more than $300 on textbooks during the spring 2020 semester, on average, the school collected data for a random sample of 76 studentss. If the calculated value for the associated test statistic equals +1.9921 and the population standard deviation is unknown, determine the p-VALUE for the associated test
Answer #1

As we are testing whether the money spent is more than $300, therefore this is a one tailed test. For n - 1 = 75 degrees of freedom, we get from the t distribution tables that:

P( t75 < 1.9921) = 0.975

Therefore, p = P( t75 < 1.9921) = 1 - 0.975 = 0.025

Therefore 0.025 is the required p-value for the associated test here.